Calculate the average speed (in kilometers per hour – hint: you will need to convert your time into units of hours) of Larry, who runs to the store 5 km away, in 15 minutes
Answer by stanbon(75887)

Calculate the average speed (in kilometers per hour – hint: you will need to convert your time into units of hours) of Larry, who runs to the store 5 km away, in 15 minutes
------
rate = 5 km/15min = [(1/3)km/min] (60 min/1 hr) = (1*60)/(3*1) = 20 km/hr
===========================
Cheers,
Stan H.